# CytoNet: A Foundation Model for the Human Cerebral Cortex

### v1

@@
CytoNet: A Foundation Model for the Human Cerebral Cortex
Christian Schiffer Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Zeynep Boztoprak Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Jan-Oliver Kropp Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Julia Thönnißen Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Katia Berr Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Hannah Spitzer Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Institute for Stroke and Dementia Research (ISD), LMU University Hospital, LMU Munich, Germany Katrin Amunts Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Timo Dickscheid Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Computer Vision, Institute for Computational Visualistics, University of Koblenz
@@
To study how the human brain works, we need to explore the organization of the cerebral cortex and its detailed cellular architecture. We introduce CytoNet, a foundation model that encodes high-resolution microscopic image patches of the cerebral cortex into highly expressive feature representations, enabling comprehensive brain analyses. CytoNet employs self-supervised learning using spatial proximity as a powerful training signal, without requiring manual labelling. The resulting features are anatomically sound and biologically relevant. They encode general aspects of cortical architecture and unique brain-specific traits. We demonstrate top-tier performance in tasks such as cortical area classification, cortical layer segmentation, cell morphology estimation, and unsupervised brain region mapping. As a foundation model, CytoNet offers a consistent framework for studying cortical microarchitecture, supporting analyses of its relationship with other structural and functional brain features, and paving the way for diverse neuroscientific investigations.
Recent advances in artificial intelligence have shown that large-scale foundation models can tackle problems once considered intractable, ranging from predicting protein structures [129] to powerful vision [131, 143] and language [139, 142] models. At the core of these successes is self-supervised learning, which extracts expressive features from massive unannotated datasets by generating implicit training signals from the data itself [122, 118]. This transition from narrow, task-specific systems to general-purpose representations marks a paradigm shift for computational problem-solving across science, technology, and industry.
@@
[…]
Here we introduce CytoNet, a foundation model for human cortical organization trained with self-supervised learning on millions of microscopic image patches from over 4,000 histological sections of ten postmortem brains. CytoNet encodes local cytoarchitectonic patterns into expressive feature representations using the proposed SpatialNCE loss, a contrastive objective that leverages anatomical proximity: patches from nearby cortical locations are treated as similar, whereas those from distant locations are treated as dissimilar. This strategy captures biologically meaningful variation without manual annotations and yields a feature space that generalizes across brains and scales. CytoNet supports applications ranging from area and layer segmentation to unsupervised clustering and comparative analysis of cortical organization. By embedding image patches into a common feature space, it enables cross-brain comparisons that reveal both shared principles of cortical architecture and brain-specific differences.
@@ 1 Results
We present the results of CytoNet from three complementary perspectives. First, we present the pretraining strategy that allowed the model to learn from large-scale histological data without annotations. Second, we analyze the learned feature space, using dimensionality reduction and clustering metrics to assess how it reflects local, global, and inter-subject cytoarchitectonic variation. Third, we demonstrate its utility in downstream applications, including prediction of structural variation, cortical area and layer segmentation, and data-driven area discovery. Together, these results demonstrate CytoNet’s scalability and versatility, showing how a single representation supports diverse analyses.
@@ 1.1 CytoNet learns cortical cytoarchitecture from spatial proximity
$\mathcal{L}_{i}=-\frac{1}{\sum_{i\neq j}\omega_{ij}}\sum_{i\neq j}\omega_{ij}\log\frac{\exp\left(z_{i}^{\top}z_{j}/\tau\right)}{\sum_{k\neq i}\exp\left(z_{i}^{\top}z_{k}/\tau\right)},$ (1)
with similarity weights
@@
All sections were co-registered to the MNI Colin 27 single subject reference space [83], enabling consistent distance computation across brains. Crucially, the model never receives spatial coordinates, and therefore similarity must be inferred from image content. Unlike augmentation-based methods that impose semantic invariance externally, SpatialNCE exploits the intrinsic continuity of brain organization, encouraging features that capture shared cytoarchitectonic properties (e.g., cell density, lamination) while suppressing confounds (e.g., staining variation or morphology).
@@ 1.2 CytoNet encodes cytoarchitectonic organization
We examined how CytoNet-ViT (1M) organizes cytoarchitectonic information in its feature space (Figure 2, top [^1] Interactive versions of selected figures are available at https://go.fzj.de/cytonet-interactive.). CytoNet-ViT (1M) is a hybrid ResNet50–ViT-B model pretrained on one million cortical patches. 2D UMAP embeddings [109] of patches from all ten brains revealed distinct brain-specific manifolds (Calinski–Harabasz index ([77], CHI) 2517.84) with consistent internal organization: atlas labels indicating different cytoarchitectonic areas [116] clustered coherently (CHI 719.66). The second UMAP dimension showed a subdivision of clusters at the central sulcus, which is an important anatomical landmark separating motor and somatosensory areas. The tenth brain (B99), which was not included in the pretraining phase, appeared more compact than the remaining nine. Yet, it showed comparable internal structure, indicating generalization beyond the training set.
To assess how well CytoNet features express cytoarchitectonic similarity across brains, pairwise cosine similarities of feature vectors were mean-aggregated into similarity matrices grouped by atlas areas (Figure 2, bottom). Resulting matrices exhibited a strong block structure, i.e., there was higher similarity within areas than between them. These patterns were highly correlated across subjects (Pearson $r=0.88\pm 0.09$), confirming that CytoNet encodes stable inter-area relationships. The tenth brain again showed elevated overall similarity and weaker block structure, consistent with the reduced specificity observed in the UMAP embeddings.
@@
Further, we studied the attention maps of the class token in the first vision transformer layer to visualize how the model attends to cytoarchitectonic structures. Figure 3 shows attention scores for example patches from the primary visual cortex hOc1 (Brodmann area 17, [88]), the (anterior) primary motor cortex area 4a [82] and area 3b of the primary somatosensory cortex [86]. The attention scores reveal how the model attends to the composition of cortical layers that define cytoarchitecture. In particular, attention heads focusing on prominent landmarks like the stripe of Gennari in hOc1, Betzt giant cells in layer V of 4a, and a pronounced layer V in 3b were identified.
@@ 1.3 Applications of CytoNet
The representations learned by CytoNet provide a versatile foundation for multiple downstream applications in human brain mapping. To demonstrate their utility, we evaluated CytoNet features in three complementary settings: (1) correlation with established structural and cytoarchitectonic properties, (2) supervised mapping tasks such as brain area classification and cortical layer segmentation, and (3) exploratory analyses for data-driven discovery of new or refined cortical areas.
@@
Abbildung 4: Comparison of predictive performance between intensity profiles and CytoNet-ViT (1M) features in B2020. A: Linear regression models using varying subsets of PCA components revealed substantially higher $R^{2}$ scores for CytoNet features compared to intensity profiles across all evaluated structural and morphological properties. Reported values reflect the average $R^{2}$ across 5-fold cross-validation. B: Absolute feature importance scores —derived from regression coefficients for the first 32 PCA components of CytoNet features— showed that components 1–3 were strongly associated with spatial location in MNI space and the density of cortical layer IV, while other properties were predominantly encoded in higher components. LI to L6 denote cortical layers I to VI. C: The cumulative explained variance across PCA components indicates that CytoNet features capture substantially more variance than intensity profiles.
@@
We assessed which anatomical properties are reflected in the feature space of CytoNet-ViT (1M), focusing on morphological and cytoarchitectonic variables extracted from B2020, the BigBrain dataset [98]. The predictive power of CytoNet features was compared to that of intensity profiles extracted from the BigBrain dataset [112], which also aim to capture local cytoarchitectonic composition. Intensity profiles mainly reflected anterior–posterior position, whereas CytoNet captured all three spatial axes as well as morphological properties such as cortical thickness, curvature, and cutting angle (Figure 4, A). It also enabled accurate prediction of cortical layer thicknesses and layer-wise cell densities, which intensity profiles from the 20 micrometer model failed to represent consistently (Figure 4, A, left columns). Across all evaluated properties, higher-dimensional PCA projections of CytoNet features generally improved predictive accuracy, while performance of intensity profile projections remained largely constant (Figure 4, B). This indicates that CytoNet encodes a richer set of structural cues that are distributed across many dimensions of the feature space, enabling fine-grained modeling of complex cytoarchitectonic patterns. Consistent with this finding, the cumulative explained variance of the PCA projections was markedly higher for CytoNet features than for intensity profiles, suggesting that CytoNet learns a more structured and informative representation space for downstream analysis (Figure 4, C).
Supervised mapping of brain areas and cortical layers
@@
Abbildung 5: Performance of cytoarchitectonic brain area classification using CytoNet. A: Macro-F1 scores obtained by linear probing of different models. Mean and standard deviation over three training runs are reported. See supplementary Table 5 for more detailed scores. If applicable, the number of pretraining samples are indicated after the model name. B: Distribution of prediction errors of CytoNet-ViT (1M) by error distance for seen, transfer, and unseen brains. Error distance was defined as the number of hops between the predicted and true brain area in the adjacency graph of the Julich Brain Atlas 3.1 [116], where 1-hop errors correspond to directly adjacent areas, and larger distances reflect increasing topological separation. C: Boxplots of the logit margins for CytoNet-ViT (1M) predictions stratified by error distance. The logit margin —the difference between the top two logits— serves as a proxy for model confidence and distance to the decision boundary [141]. Correct predictions reveal higher confidence, while incorrect predictions show decreasing confidence with increasing error distance.
@@
In area classification, CytoNet consistently outperformed models trained from scratch, SimCLR [118], and supervised contrastive baselines [132] across seen, transfer, and unseen brains (Figure 5, A and supplementary Table 5, see supplementary Section 3.5 for model details). CytoNet-ViT (1M) reached the best scores on seen brains (macro-F1 0.69 with linear probing, 0.71 with finetuning), surpassing all supervised alternatives and SimCLR. On the transfer brain, where no area annotations were available for training, CytoNet again achieved the highest scores. Even on the unseen brain, which was excluded from both pretraining and classifier training, CytoNet models remained competitive, with compact CNN-based variants showing the strongest generalization. In contrast, SimCLR variants performed poorly across all splits, often below training from scratch. Retrieval analysis confirmed this shortcut learning, showing that SimCLR features clustered images by tissue morphology and vascular patterns rather than area identity (see supplementary Figure 10).
@@
Error analysis (Figure 5, B) revealed that CytoNet’s misclassifications were largely confined to borders between adjacent areas: $\sim$80% of errors in seen and transfer brains occurred within 1 hop of the correct area in the atlas adjacency graph, and over 95% within 2 hops. For the unseen brain, these proportions were somewhat lower (52% and 85%), but errors remained topologically plausible. Confidence, measured by logit margin [141], was significantly higher for correct predictions and decreased systematically with error distance (Figure 5, F). These findings indicate that CytoNet’s errors mirror the uncertainties faced by human experts, with most mistakes arising at difficult borders rather than random misclassifications.
@@
Abbildung 6: Results for cortical layer segmentation and data-driven discovery of brain areas. A: Macro-F1 scores from linear probing of different models with varying amounts of training data (mean ± SD across five-fold cross-validation; evaluated on a dedicated test set, see supplementary Table 7). If applicable, the number of pretraining samples are indicated after the model name. B: Example segmentation of cortical layers in area 4a (primary motor cortex, [82]) for scratch (200k) and CytoNet-ViT (1M) models for increasing fractions of the training set. C: K-means clusters (indicated by colors) of CytoNet-ViT (1M) features in the vicinity of frontal pole areas Fp1 and Fp2 in brain B66, with clusters pre-localized using Julich Brain Atlas 3.1 probability maps (threshold 50%) and annotations. Hemispheres are shown separately for visualization, but clustering was performed jointly.
@@
Cortical layer segmentation further demonstrated the richness and data efficiency of CytoNet features. Using only 1% of the annotated dataset (7 training patches), linear probing of CytoNet-ViT (1M) reached a macro-F1 of 0.63, i.e., over four times the scratch baseline (0.15) and well above SimCLR-pretrained models (0.49). With 5% of training data, CytoNet already achieved macro-F1 of 0.74, while baselines required 20% of data to reach similar performance (Figure 6, A, B and supplementary Table 7). Finetuning CytoNet improved accuracy with larger training datasets but caused overfitting in small-data regimes and instability in transformer backbones, whereas linear probing remained robust in both cases.
Data-driven discovery of brain areas
@@
We finally asked whether CytoNet features can support exploratory, data-driven refinement of brain parcellations through clustering. Accurately identifying and delineating brain areas remains a central challenge in brain mapping and analysis, particularly in regions lacking anatomical landmarks. While historical atlases such as Brodmann’s cytoarchitectonic map [73] provide a foundational parcellation, many regions have since been identified based on refined histological evidence, often by subdividing existing areas. As a case study, we examined the frontal pole, which was initially defined as Brodmann area 10, but later subdivided into Fp1 and Fp2 by [100]. We treated the region as if no subdivision were known and tested whether clustering in CytoNet’s feature space could recover the distinction.
@@
CytoNet features from brain B66 were clustered using $k$-means, applied either to points pre-localized by probabilistic maps or to points annotated as Fp1/Fp2 (Figure 6, C). In the probabilistic setting, one cluster aligned well with Fp2, while Fp1 points split into multiple clusters with inter-hemispheric differences. This resulted in high purity with respect to anatomical labels ($0.97\pm 0.02$) and hemisphere ($0.91\pm 0.02$). In the annotation-based setting, clustering robustly separated Fp1 and Fp2, with mean purity $0.94\pm 0.001$. Here, hemisphere information was more evenly distributed across clusters (purity $0.52\pm 0.001$), indicating that areal identity dominated. Manually identifying cluster 0 with Fp1 and cluster 1 with Fp2 results (supplementary Figure 9) yielded an accuracy of $94.75\%$.

### v2

@@
CytoNet: A Foundation Model for the Human Cerebral Cortex at Cellular Resolution
Christian Schiffer Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Zeynep Boztoprak Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Jan-Oliver Kropp Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Julia Thönnißen Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Katia Berr Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Hannah Spitzer Institute of Computational Biology, Computational Health Center, Helmholtz Munich, Munich, Germany Institute for Stroke and Dementia Research (ISD), LMU University Hospital, LMU Munich, Germany Katrin Amunts Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Cécile & Oscar Vogt Institute for Brain Research, University Hospital Düsseldorf, Düsseldorf, Germany Timo Dickscheid Institute of Neuroscience and Medicine (INM-1), Research Centre Jülich, Jülich, Germany Helmholtz AI, Research Centre Jülich, Jülich, Germany Computer Vision, Institute for Computational Visualistics, University of Koblenz, Koblenz, Germany
@@
Studying the cellular architecture of the human cerebral cortex is critical for understanding brain organization and function. It requires investigating complex texture patterns in histological images, yet automatic methods that scale across whole brains are still lacking. Here we introduce CytoNet, a foundation model trained on 1 million unlabeled microscopic image patches from over 4,000 histological sections spanning ten postmortem human brains. Using co-localization in the cortical sheet for self-supervision, CytoNet encodes complex cellular patterns into expressive and anatomically meaningful feature representations. CytoNet supports multiple downstream applications, including area classification, laminar segmentation, quantification of microarchitectural variation, and data-driven mapping of previously uncharted areas. In addition, CytoNet captures microarchitectural signatures of macroscale functional organization, enabling decoding of functional network parcellations from cytoarchitectonic features. Together, these results establish CytoNet as a unified framework for scalable analysis of cortical microarchitecture and for linking cellular architecture to structure–function organization in the human cerebral cortex.
1 Introduction
@@
Recent advances in artificial intelligence (AI) suggest a route to address this challenge through large-scale foundation models that learn general-purpose representations from data [12]. Such models have enabled breakthroughs ranging from protein structure prediction [45] to vision [69, 62] and language [21, 61]. A key driver of these successes is self-supervised learning, which extracts expressive features from massive unannotated datasets by generating implicit training signals from the data itself [39, 19]. In this novel paradigm, computational approaches shift from designing specialized analysis pipelines to learning general feature representations that directly support data-driven investigation, prediction, and discovery in large-scale datasets. In neuroimaging, analogous approaches are emerging to learn transferable representations that enable adaptation to diverse downstream prediction tasks, for example from MRI [77] and fMRI [28] datasets. At the cellular scale, histological data provides direct access to cortical microarchitecture, motivating foundation models that learn transferable cytoarchitectonic representations from whole-brain microscopy and relate them to macroscale cortical organization and functional systems.
To this end, we introduce CytoNet, a foundation model for human cortical organization, trained with self-supervised learning on 1 million microscopic image patches from over 4,000 cell-body–stained histological sections spanning ten postmortem human brains. CytoNet provides a data-driven approach to quantifying cortical cytoarchitecture by embedding local histological patterns into a continuous feature space that transfers across brains and downstream tasks. Trained using the proposed SpatialNCE loss, a contrastive learning objective that leverages spatial co-localization within the cortical sheet as a self-supervised training signal, CytoNet learns to encode image patches into expressive representations of cytoarchitecture, without requiring manual annotations. We show that these representations are anatomically meaningful and robust. They generalize across brains and spatial scales, while capturing both common cytoarchitectonic structure and brain-specific variation that facilitate cross-brain comparison at whole-brain scale. This shared representation enables cytoarchitectonic mapping, prediction, and exploration at scale, beyond what is practical with task-specific approaches. In cytoarchitectonic area mapping, CytoNet enables accurate classification of over 100 cortical areas, outperforming training from scratch, general self-supervised baselines [19], and specialized approaches [72]. CytoNet also supports label-efficient cortical layer segmentation, producing reliable laminar maps with only a handful of annotated training samples. Beyond mapping, CytoNet features accurately model structural variables, including layer-wise cell densities, outperforming established profile-based descriptors [84]. If cytoarchitecture provides a substrate for systems-level specialization [59, 41], microarchitectural representations should carry information about macroscale functional organization. We therefore test whether CytoNet features can decode widely used functional network parcellations, using spatially aware validation to control for spatial smoothness. Finally, we provide a proof-of-concept that CytoNet features can drive unsupervised mapping of previously unknown regional structures. Together, these results establish CytoNet as a scalable framework for learning, quantifying, and comparing cortical cytoarchitecture from whole-brain histology, enabling systematic analyses across thousands of sections and entire human brains.
@@ 2 Results
We present the results of CytoNet from three complementary perspectives. First, we introduce a pretraining strategy that allowed the model to learn from large-scale histological data without annotations. Second, we analyze the learned feature space, using dimensionality reduction and clustering metrics to assess how it reflects local, global, and inter-subject cytoarchitectonic variation. Third, we demonstrate its utility in downstream applications, including prediction of structural variations, cortical area and layer segmentation, decoding of functional organization, and unsupervised mapping of uncharted areas. Together, these results demonstrate CytoNet’s scalability and versatility, showing how a single representation supports diverse analyses.
@@ 2.1 CytoNet learns cortical cytoarchitecture from spatial proximity
$\mathcal{L}_{i}=-\frac{1}{\sum_{j\neq i}\omega_{ij}}\sum_{j\neq i}\omega_{ij}\log\frac{\exp\left(z_{i}^{\top}z_{j}/\tau\right)}{\sum_{k\neq i}\exp\left(z_{i}^{\top}z_{k}/\tau\right)},$ (1)
with temperature $\tau$ and similarity weights defined by the radial basis function (RBF) kernel
@@
with bandwidth $\sigma$. All sections were co-registered to the MNI Colin 27 single subject reference space [43], enabling consistent distance computation across brains. Crucially, the model never receives spatial coordinates, and therefore similarity must be inferred from image content. Unlike augmentation-based methods that impose semantic invariance externally, SpatialNCE exploits the intrinsic continuity of brain organization, encouraging features that capture shared cytoarchitectonic properties (e.g., cell density, lamination) while suppressing confounds (e.g., staining variation or morphology).
@@ 2.2 CytoNet encodes cytoarchitectonic organization
We examined how CytoNet-ViT (1M) organizes cytoarchitectonic information in its feature space (Fig. 2, top [^1] Interactive versions of selected figures are available at https://go.fzj.de/cytonet-interactive.). CytoNet-ViT (1M) is a hybrid ResNet50–ViT-B model pretrained on one million cortical patches. 2D UMAP embeddings [57] of patches from all ten brains revealed distinct brain-specific manifolds (Calinski–Harabasz index ([14], CHI) 2517.84) with consistent internal organization: atlas labels indicating different cytoarchitectonic areas [6] clustered coherently (CHI 719.66). The second UMAP dimension showed a subdivision of clusters at the central sulcus, which is an important anatomical landmark separating motor and somatosensory areas. The tenth brain (B99), which was not included in the pretraining phase, appeared more compact than the remaining nine. Yet, it showed comparable internal structure, indicating generalization beyond the training set.
To assess how well CytoNet features express cytoarchitectonic similarity across brains, pairwise cosine similarities of feature vectors were mean-aggregated into similarity matrices grouped by atlas areas (Fig. 2, bottom). Resulting matrices exhibited a strong block structure, i.e., there was higher similarity within areas than between them. These patterns were highly correlated across subjects (Pearson $r=0.88\pm 0.09$), confirming that CytoNet encodes stable inter-area relationships. The tenth brain again showed elevated overall similarity and weaker block structure, consistent with the reduced specificity observed in the UMAP embeddings.
@@
Further, we studied the attention maps of the class token in the first vision transformer layer to visualize how the model attends to cytoarchitectonic structures. Fig. 3 shows attention scores for example patches from the primary visual cortex hOc1 (Brodmann area 17, [5]), the (anterior) primary motor cortex area 4a [34] and area 3b of the primary somatosensory cortex [35]. The attention scores reveal how the model attends to the composition of cortical layers that define cytoarchitecture. In particular, attention heads focusing on prominent landmarks like the stripe of Gennari in hOc1, Betzt giant cells in layer V of 4a, and a pronounced layer V in 3b were identified.
@@ 2.3 Applications of CytoNet
The representations learned by CytoNet provide a versatile foundation for multiple downstream applications in human brain mapping. To demonstrate their utility, we evaluated CytoNet features in four complementary settings: (1) correlation with established structural and cytoarchitectonic properties, (2) decoding of macroscale functional organization, (3) supervised mapping tasks such as brain area classification and cortical layer segmentation, and (4) analyses for data-driven mapping of new or refined cortical areas.
@@
Abbildung 4: Decoding structural properties and functional parcellations from CytoNet representations in B2020. A: Linear regression models using varying subsets of PCA components revealed substantially higher $R^{2}$ scores for CytoNet features compared to intensity profiles across all evaluated structural and morphological properties. Reported values reflect the $R^{2}$ across 5-fold cross-validation. CytoNet PCA-NN indicates the first NN PCA components of CytoNet features. B: Absolute feature importance scores derived from regression coefficients for the first 32 PCA components of CytoNet features showed that components 1–3 were strongly associated with spatial location in MNI space and the density of cortical layer IV, while other properties were predominantly encoded in higher components. LI to L6 denote cortical layers I to VI. C: The cumulative explained variance across PCA components indicates that CytoNet features capture substantially more variance than intensity profiles. D: Decoding functional parcellation labels (Yeo7, Yeo17; [87], and DiFuMo64; [23]) from different feature sets using multinomial logistic regression. Reported values reflect macro-F1 across 5-fold cross-validation. CytoNet features outperformed histological (Hist-G, [65]) and microstructural (Micro-G, [71]) gradient baselines as well as a low-dimensional functional-gradient reference (Func-G, [53]). A baseline using only fs_LR vertex coordinates quantifies the contribution of cortical location. Concatenating CytoNet features with coordinates (CytoNet PCA-32+C) yields the highest score, indicating complementary information beyond spatial position.
@@
We assessed which anatomical properties are reflected in the feature space of CytoNet-ViT (1M), focusing on morphological and cytoarchitectonic variables extracted from B2020, the BigBrain dataset [3]. The predictive power of CytoNet features was compared to that of intensity profiles extracted from the BigBrain dataset [86], which also aim to capture local cytoarchitectonic composition. Intensity profiles mainly reflected anterior–posterior position, whereas CytoNet captured all three spatial axes as well as morphological properties such as cortical thickness, curvature, and cutting angle (Fig. 4, A). It also enabled accurate prediction of cortical layer thicknesses and layer-wise cell densities, which intensity profiles from the 20 micrometer model failed to represent consistently (Fig. 4, A, left columns). Across all evaluated properties, higher-dimensional PCA projections of CytoNet features generally improved predictive accuracy, while performance of intensity profile projections remained largely constant (Fig. 4, B). This indicates that CytoNet encodes a richer set of structural cues that are distributed across many dimensions of the feature space, enabling fine-grained modeling of complex cytoarchitectonic patterns. Consistent with this finding, the cumulative explained variance of the PCA projections was markedly higher for CytoNet features than for intensity profiles, suggesting that CytoNet learns a more structured and informative representation space for downstream analysis (Fig. 4, C).
Decoding functional organization from cytoarchitecture
We evaluated whether CytoNet representations can decode macroscale functional organization. To this end, we trained multinomial logistic regression models to predict network labels from the Yeo7 and Yeo17 parcellations [87] as well as the DiFuMo atlas [23] with 64 functional modes, denoted as DiFuMo64 (Fig. 4, D).
Cortical location alone provided a strong baseline, consistent with the spatial smoothness of functional networks. Using only fs_LR (32k) vertex coordinates [81] achieved macro-F1 of $0.40\pm 0.03$ (Yeo7), $0.28\pm 0.03$ (Yeo17), and $0.25\pm 0.03$ (DiFuMo64). CytoNet features provided substantial additional predictive signal, and performance increased systematically with feature dimensionality: PCA–1 reached $0.21\pm 0.02$, $0.10\pm 0.01$, and $0.04\pm 0.01$ (Yeo7/Yeo17/DiFuMo64), PCA1–2 reached $0.33\pm 0.04$, $0.18\pm 0.03$, and $0.12\pm 0.01$, and PCA–32 reached $0.63\pm 0.03$, $0.51\pm 0.03$, and $0.31\pm 0.02$. To test whether CytoNet’s advantage over coordinates could be explained by inherent spatial smoothness alone, we performed a spin-based spatial null on the parcellation labels in fs_LR space and assessed the significance of the CytoNet–coordinate macro-F1 delta. The improvement was significant for Yeo17 (10,000 spins; one-sided $p=0.0036$; Holm–Bonferroni corrected across the three parcellations tested), Yeo7 showed an uncorrected effect ($p=0.042$); DiFuMo64 was not significant ($p=0.98$). Notably, while overall decoding was higher for the coarser Yeo7 parcellation, evidence for an improvement beyond spatial smoothness was stronger for the finer-grained Yeo17 labels under the spin null, consistent with CytoNet capturing microarchitectural information beyond coarse spatial gradients.
Compared to gradient baselines, CytoNet PCA–32 outperformed both histological (Hist-G, [65]) and microstructural (Micro-G, [71]) gradients across all targets (Hist-G1+2 $21\pm 0.02, 0.0$$9\pm 0.01, 0.0$$03\pm 0.01$; Micro-G1+2 $0.34\pm 0.04$, $17\pm 0.02, 0.0$$06\pm 0.01$). Interestingly, it also exceeded a low-dimensional functional-gradient reference (Func-G1+2: $0.51\pm 0.02$, $0.40\pm 0.03$, $0.17\pm 0.02$; [53]), indicating that discrete network assignments can require feature combinations not fully captured by only two functional gradient axes.
Finally, combining CytoNet with explicit spatial information yielded the best performance. Concatenating vertex coordinates with CytoNet PCA–32 increased macro-F1 to $0.67\pm 0.03$ (Yeo7), $0.56\pm 0.03$ (Yeo17), and $0.44\pm 0.02$ (DiFuMo64), suggesting that cytoarchitectonic cues and cortical position provide complementary information for decoding functional parcellations.
Mapping of brain areas and cortical layers
@@
Abbildung 5: Performance of cytoarchitectonic brain area classification using CytoNet. A: Macro-F1 scores obtained by linear probing of different models across three training runs. See supplementary Table 5 for more detailed scores. If applicable, the number of pretraining samples are indicated after the model name. B: Distribution of prediction errors of CytoNet-ViT (1M) by error distance for seen, transfer, and unseen brains. Error distance was defined as the number of hops between the predicted and true brain area in the adjacency graph of the Julich Brain Atlas 3.1 [6], where 1-hop errors correspond to directly adjacent areas, and larger distances reflect increasing topological separation. C: Boxplots of the logit margins for CytoNet-ViT (1M) predictions stratified by error distance. The logit margin —the difference between the top two logits— serves as a proxy for model confidence and distance to the decision boundary [60]. Correct predictions reveal higher confidence, while incorrect predictions show decreasing confidence with increasing error distance.
@@
In area classification, CytoNet consistently outperformed models trained from scratch, SimCLR [19], and supervised contrastive baselines [72] across seen, transfer, and unseen brains (Fig. 5, A and supplementary Table 5, see Section 4.5 for model details). CytoNet-ViT (1M) reached the best scores on seen brains (macro-F1 0.69 with linear probing, 0.71 with finetuning), surpassing all supervised alternatives and SimCLR. On the transfer brain, where no area annotations were available for training, CytoNet again achieved the highest scores. Even on the unseen brain, which was excluded from both pretraining and classifier training, CytoNet models remained competitive, with compact CNN-based variants showing the strongest generalization. In contrast, SimCLR variants performed poorly across all splits, often below training from scratch. Retrieval analysis confirmed this shortcut learning, showing that SimCLR features clustered images by tissue morphology and vascular patterns rather than area identity (see supplementary Fig. 10).
@@
Error analysis (Fig. 5, B) revealed that CytoNet’s misclassifications were largely confined to borders between adjacent areas: $\sim$80% of errors in seen and transfer brains occurred within 1 hop of the correct area in the atlas adjacency graph, and over 95% within 2 hops. For the unseen brain, these proportions were somewhat lower (52% and 85%), but errors remained topologically plausible. Confidence, measured by logit margin [60], was significantly higher for correct predictions and decreased systematically with error distance (Fig. 5, F). These findings indicate that CytoNet’s errors mirror the uncertainties faced by human experts, with most mistakes arising at difficult borders rather than random misclassifications.
@@
Abbildung 6: Results for cortical layer segmentation and data-driven mapping of brain areas. A: Macro-F1 scores from linear probing of different models with varying amounts of training data across five-fold cross-validation, evaluated on a dedicated test set (also see supplementary Table 7). If applicable, the number of pretraining samples are indicated after the model name. B: Example segmentation of cortical layers in area 44 ([7]) from the test set for scratch and CytoNet-ViT (1M) models for increasing fractions of the training set. C: K-means clusters of CytoNet-ViT (1M) features in the vicinity of frontal pole areas Fp1 and Fp2 in brain B66, with clusters pre-localized using Julich Brain Atlas 3.1 probability maps (threshold 50%) and annotations. Clustering was performed separately based on probabilities (left) and annotations (right). Colors indicate cluster assignment. Results are shown per hemisphere, but clustering was performed jointly across hemispheres.
@@
Cortical layer segmentation further demonstrated the richness and data efficiency of CytoNet features. Using only 1% of the annotated dataset (7 training patches), linear probing of CytoNet-ViT (1M) reached a macro-F1 of 0.63, i.e., over four times the scratch baseline (0.15) and well above SimCLR-pretrained models (0.49). With 5% of training data, CytoNet already achieved macro-F1 of 0.74, while baselines required 20% of data to reach similar performance (Fig. 6, A, B and supplementary Table 7). Finetuning CytoNet improved accuracy with larger training datasets but caused overfitting in small-data regimes and instability in transformer backbones, whereas linear probing remained robust in both cases.
Data-driven mapping of brain areas
@@
We finally asked whether CytoNet features can support exploratory, data-driven refinement of brain parcellations through clustering. Accurately identifying and delineating brain areas remains a central challenge in brain mapping and analysis, particularly in regions lacking anatomical landmarks. While historical cytoarchitectonic map (e.g., [13]) provide a foundational parcellation, many regions have since been identified based on refined histological evidence, often by subdividing existing areas. As a case study, we examined the frontal pole, which was initially defined as Brodmann area 10, but later subdivided into Fp1 and Fp2 by [11]. We treated the region as if no subdivision were known and tested whether clustering in CytoNet’s feature space could recover the distinction.
@@
CytoNet features from brain B66 were clustered using $k$-means, applied either to points pre-localized by probabilistic maps or to points annotated as Fp1/Fp2 (Fig. 6, C). In the probabilistic setting, one cluster aligned well with Fp2, while Fp1 points split into multiple clusters with inter-hemispheric differences. This resulted in high purity with respect to anatomical labels ($0.97\pm 0.02$) and hemisphere ($0.91\pm 0.02$). In the annotation-based setting, clustering robustly separated Fp1 and Fp2, with mean purity $0.94\pm 0.001$. Here, hemisphere information was more evenly distributed across clusters (purity $0.52\pm 0.001$), indicating that areal identity dominated. Manually identifying cluster 0 with Fp1 and cluster 1 with Fp2 (supplementary Fig. 9) yielded an accuracy of $94.75\%$.